A ox ค. eing from a hunter encounters a 600 m tal fence and attempts to ump it. The ox umps with an initial velocity of 7.05 m/s at an angle or 45 0.. be from the fence. By how much does the fox clear the fence? Treat the fox as a particle. ning the ump 1.02 m
Answer #1

Solution-

Let us ignore air resistance
assuming fence sits on level ground
assume up = positive direction and ground level is the origin

horizontal velocity component= 7.05cos45 =4.98 m/s

To reach the fence from launch point the time needed
1.92 m / 4.98 m/s = 0.38 s

vertical velocity component 7.05sin45 = 4.98 m/s

vertical position when he reaches the fence

y = y₀ + v₀t + ½gt^2

y = 0 + 4.98(0.38) + ½(-9.81)0.38^2
y = 1.18 m

fence clearance= 1.18 - 0.690

=0.49 m